# Meritocracy and Its Discontents: Long-run Effects of Repeated School Admission Reforms

### v1

@@ A.4.2 Data on the National Production of Career Elites
In the second set of long-run analysis, we focus on a specific group of career elites, i.e., higher civil officials for whom we have complete count data.
@@
Our main data source is Hata (1981), which provides not only the list of all individuals who passed the Higher Civil Service Examinations (HCSE, Bunkan Koutou Shiken in Japanese), but also their biographical information available as of 1981 (Hata, 1981, Section 3). The HCSE were highly selective national qualification exams held annually from 1894 to 1947. The administrative division of the HCSE consisted of two parts: the preliminary exams and the main exams. Until 1922, graduates of the law departments of imperial universities were exempted from the preliminary exams, but had to take the main exams (Hata, 1981, pp.663-666; Spaulding, 1967, Chapter 12).
@@
To proceed to an empirical analysis, we redefine cohort by “the year of entering a higher school or its equivalent” estimated above (so that we can see which cohorts are exposed to Capp as opposed to Dapp). That is, to create cohort-level data, we count the number of exam passers (i.e., those who passed the administrative division of the HCSE) and the number of top-ranking officials by the year in which they entered a higher school or its equivalent. Our dataset consists of 6,255 individuals who entered a higher school or its equivalent in 1898–1930 and passed the HCSE in 1901–1941.
Out of 6,255 exam passers, 3,490 individuals or 55.8% are Schools 1–8 graduates, 4,767 individuals or 76.2% are Imperial University graduates, and 982 individuals or 15.7% are top-ranking officials. The share of Imperial University graduates is greater than that of Schools 1–8 graduates because the number of higher schools increased from 8 to 25 in 1919–1930. Among 982 top-ranking officials, 701 officials or 71.4% are Schools 1–8 graduates, and 891 officials or 90.7% are Imperial University graduates. The HCSE consisted of preliminary and main exams, and because Imperial University law graduates were the only group exempted from the administrative preliminary exams from 1894 to 1922, they had substantial advantages in passing the HCSE (Hata, 1981, pp.663-666; Spaulding, 1967, Chapter 12). This exemption, however, does not explain why Imperial University graduates had such a high share in the top-ranking officials.
@@ A.4.3 Variation in Years Taken to be Promoted to Top-ranking Positions
One potential threat to our identification in Table 4 Panel (b) is a possibility that the number of available top-ranking positions happened to have increased during the periods of Capp. However, we argue that even if this was the case, it is not likely to affect our results, since our cohort is defined by the year of entering a higher school or its equivalent, not by the year of becoming top-ranking officials. Namely, as long as individuals in a given cohort were not promoted to a top-ranking position in the same year, a potential correlation between the number of top-ranking positions and the lagged periods of centralized admissions does not bias our results.
To examine this possibility, we first randomly selected two cohorts exposed to Capp (1903 and 1916) and two cohorts exposed to Dapp (1913 and 1922). For all top-ranking officials who graduated from Schools 1–8 in these cohorts, we searched the years in which these officials were appointed to their first top-ranking positions. Using online searches to find biographical information for each official, we obtained necessary information for 82% of these officials. As shown in Appendix Figure A.9, within each cohort, the number of years taken from entering a higher school to the appointment for the first top-ranking position varied widely from 20 to 30 years. This within-cohort variation is a sum of two variations, the first variation in the number of years taken from entering a higher school to passing the exams and the second variation in the number of years taken from passing the exams to becoming a top-ranking official. Therefore, it is unlikely that the result in Table 4 Panel (b) is driven by a greater number of available top-ranking positions that coincided with the periods of Capp.
@@ A.4.4 Connections and Signaling Hypotheses
We investigate whether connections or signaling was an important mechanism through which the meritocratic centralization increased the total number of top-ranking officials. First, to explore the connection channel, we reexamine the result of Table 4 Panel (b) Column (3) by controlling for the number of exam passers from the same schools (Schools 2–8) in the same cohort, instead of the total number of exam passers. This variable takes into account the size of same-cohort colleagues from the same schools who are likely to have network values. The coefficient of Capp remains sizable (2.72) and statistically significant at the 1% level, implying 24% increase in the number of top-ranking officials from these schools relative to the mean during Dapp (11.22). Second, to investigate the signaling channel, we further restrict our sample to top-ranking officials who graduated from both Schools 2–8 and Tokyo Imperial University, the most prestigious institution with the strongest signaling value. We regress the number of this subgroup of top-ranking officials on Capp, controlling for the number of exam passers from the same schools in the same cohort. If signaling is a major channel, we expect the effect of Capp among this subgroup to be larger. The coefficient of Capp decreases to 1.76, implying 18% increase in the number of top-ranking officials from these schools relative to the mean during Dapp (10.00). This effect size is smaller than the effect size of 24% above.
@@
Notes: This figure summarizes each country and territory’s college admission system today. Dark red color (e.g. Norway): Regionally- or nationally-centralized college admissions where a single-application, single-offer assignment algorithm (well-defined rule) is used to make admissions to both public and private universities. Medium orange color (e.g. Brazil): Semi-centralized college admissions defined as either (1) there is a centralized system, but not all universities (such as private universities) are included in the single-application, single-offer system or (2) students submit a single application and receive multiple offers. Light orange color (e.g. US): Decentralized college admissions where each college defines its own admission standards and rules. Yellow with diagonal lines (e.g. Chad): Not enough information available or if the country or territory does not have tertiary institutions. We summarize the information sources at https://www.scribd.com/document/437545135/Online-Appendix191018. See Section 2 for discussions about this figure.
@@
Notes: This figure shows the locations of Schools 1–8 and the location of the Tokyo area (in dark gray color) defined as a set of 7 prefectures that are within 100 km from Tokyo (i.e., Tokyo, Chiba, Kanagawa, Saitama, Ibaraki, Tochigi, and Gunma prefectures). See Sections 2 and 3.4 for discussions about this figure.
@@
Notes: This table shows changes in the number and the admission system of Schools 1-8 from 1900 to 1930. In 1901, all schools held their exams on the same date, with the exception of newly established School 7. In 1908, all schools held their exams on the same date, with the exception of Schools 7 and 8. See Section 2.1 and Appendix Section A.1 for discussions about this table.
@@
Notes: This figure is a reprint of the assignment algorithm of the centralized admission system stated in the Ordinance of the Ministry of Education No.4 published in Government Gazette No.1419, pp.580-581, on April 27, 1917. See Sections 2 and 3.3 for an English translation and discussions.
@@
No. of public middle-school graduates 415.38 350.01 299 1410
No. of private middle-school graduates 118.49 435.12 0 1410
@@
No. of corporate executives with a positive amount of tax payment 7.99 8.91 5 705
@@ B. Application Distance
Notes: This table uses the prefecture-year level data in 1900–1930. The dependent variable is the number of students from birth prefecture $p$ who entered one of Schools 1–8 in year $t$. “Population in prefecture” is population in prefecture $p$ in year $t$. “GDP per capita in prefecture” is real gross value-added per capita in prefecture $p$ in year $t$. “Middle-school graduates in prefecture” is the number of students who graduated from middle schools in prefecture $p$ in year $t$. “Middle-school graduates in nearby prefectures” is the number of students who graduated from middle schools in the prefectures within 100 km from prefecture $p$ (excluding prefecture $p$) in year $t$. “Share of applicants to School 1 (under decentralization)” is the share of applicants to School 1 among all applicants to Schools 1–8 in prefecture $p$ under the decentralized system in 1916 (the only year for which the data is available). All variables interacted with “Centralized” are standardized to be mean 0 and standard deviation 1. We control for year fixed effects, prefecture fixed effects, and the number of higher schools other than Schools 1–8 in prefecture $p$ in year $t$. We also control for “Born in school’s prefecture,” “Born near school’s prefecture (1-100 km),” and “Born near school’s prefecture (100-300 km)” as in Table 1. Parentheses contain p-values based on standard errors clustered at the prefecture level. Square brackets contain p-values based on standard errors clustered at the year level. ***, **, and * mean significance at the 1%, 5%, and 10% levels, respectively. See Section 3.4 for discussions about this table.
@@
No. of middle-school graduates 0.00493 -0.000942 -0.150 0.00134 0.0656 -0.429 0.00194 1.345 0.00147
@@
Mean dep var 23.41 2.007 0.180 10.78 0.217 19.03 13.56 0.291 228.9 0.174
Notes: Columns 1–7 test if important institutional variables are correlated with the timing of centralization using year-level data. Columns 8–10 examine the robustness of our main short-run outcomes using year-level data. All numbers are at the national-level from 1900 to 1930. The numbers of middle school graduates, entrants, and applicants are denominated by 1,000. In all regressions, quadratic time trends (i.e. trend and trend squared, where the trend is defined by “year - 1899”) are controlled. Parentheses contain p-values based on Newey-West standard errors with the maximum lag order of 3. See Section 3.5 for discussions about this table.
Table A.8: Definitions of Career Elites in the PIR
Notes: This table provides the precise definitions of career elites, defined as a subset of individuals listed in PIR (1939). See Section 4.1 for the analysis using these variables.
@@
Notes: This figure plots the sampling rate of the high income earners listed in PIR (1939) by the income level expressed as a top percentile of the national income distribution. The sampling rates and the top income percentiles are computed from the complete count data in National Tax Bureau Yearbook. The vertical lines indicate the top 0.05% and 0.01% thresholds used in our analysis. See Section 4.1 for discussions about this figure.
@@
Notes: These figures compare the number of high income earners in each prefecture listed in PIR (1939) and the complete count of high income earners in each prefecture reported in National Tax Bureau Yearbook (1936). The vertical axis is the log of the number of individuals in the PIR who earned more than 50,000 JPY taxable income (corresponding to the top 0.01% income group) or 16,950 JPY taxable income (corresponding to the top 0.05% income group) in 1938. The horizontal axis is the log of the number of individuals in tax statistics who earned more than 30,000 JPY taxable income (corresponding to the top 0.013% income group) or 10,000 JPY taxable income (corresponding to the top 0.08% income group) in 1936 (the closest year to 1938 for which prefecture-level tax statistics are available). See Section 4.1 for discussions about these figures.
@@
Figure A.7: Long-run Impacts of Centralization: Geographical Origins of Other Career Elites
Notes: This figure shows additional difference-in-differences plots that compare the average number of elites born in prefectures inside and outside the Tokyo area by cohort (see Figure 3). The plots are based on the data from PIR (1939), which covers cohorts who were born in 1879–1894 and turned age 17 (main application age) in 1896–1911. The vertical axis shows the number of indicated elites who were born in the indicated area in the indicated cohort. The cohorts are colored according to their intensity of exposure to the centralized admissions in 1902–07, where the darker color indicates the higher intensity. See Section 4.1 for discussions about this figure.
Table A.10: Long-run Impacts: Other Elite Professionals
@@
Age 17 under centralization 0.87 0.66 0.46 0.18
$\times$ Tokyo area ($<$100 km) (0.026)** (0.048)** (0.038)** (0.420)
[0.094]* [0.125] [0.091]* [0.544]
@@
Mean dep var 3.61 2.46 1.92 1.11
Mean dep var (Tokyo area 4.76 3.13 2.52 1.16
@@
Imperial Top 0.01% Top 0.05% Medal Corporate Top Imperial
Univ. income income recipients executives politicians & Univ.
grads earners earners bureaucrats professors
A. Baseline Specification
Age 17 under centralization 3.05 0.68 1.65 2.61 1.89 0.68 0.45
  $\times$ Tokyo area ($<$100 km) (0.003)*** (0.007)*** (0.048)** (0.007)*** (0.030)** (0.019)** (0.067)*
[0.000]*** [0.026]** [0.011]** [0.000]*** [0.066]* [0.023]** [0.189]
B. Adding Control Variables
Age 17 under centralization 1.70 0.58 1.53 2.18 1.60 0.45 0.41
   $\times$ Tokyo area ($<$100 km) (0.021)** (0.013)** (0.007)*** (0.000)*** (0.014)** (0.028)** (0.095)*
[0.017]** [0.008]*** [0.020]** [0.006]*** [0.069]* [0.086]* [0.186]
Observations 611 611 611 611 611 611 611
Birth cohort FE, Birth pref. FE Yes Yes Yes Yes Yes Yes Yes
Mean dep var 8.68 1.23 5.16 6.08 7.96 1.34 0.83
Mean dep var (Tokyo area 10.38 1.41 6.57 6.61 10.91 1.21 1.02
@@ B. Adding Control Variables
Imperial Top 0.01% Top 0.05% Medal Corporate Top Imperial
Univ. income income recipients executives politicians & Univ.
grads earners earners bureaucrats professors

### v2

@@ A.4.2 Data on the National Production of Top-Ranking Higher Civil Officials
In the second set of long-run analysis, we focus on a specific group of elites, i.e., higher civil officials for whom we have complete count data.
@@
Our main data source is Hata (1981), which provides not only the list of all individuals who passed the Higher Civil Service Examinations (HCSE, Bunkan Koutou Shiken in Japanese), but also their biographical information available as of 1981 (Hata, 1981, Section 3). The HCSE were selective national qualification exams held annually from 1894 to 1947. The administrative division of the HCSE consisted of two parts: the preliminary exams and the main exams. Until 1922, graduates of the law departments of imperial universities were exempted from the preliminary exams, but had to take the main exams (Hata, 1981, pp.663-666; Spaulding, 1967, Chapter 12).
@@
To proceed to an empirical analysis, we redefine cohort by “the year of entering a higher school or its equivalent” estimated above (so that we can see which cohorts are exposed to centralization as opposed to decentralization). That is, to create cohort-level data, we count the number of exam passers (i.e., those who passed the administrative division of the HCSE) and the number of top-ranking officials by the year in which they entered a higher school or its equivalent. Our dataset consists of 6,255 individuals who entered a higher school or its equivalent in 1898–1930 and passed the HCSE in 1901–1941.
Out of 6,255 exam passers, 3,490 individuals (55.8%) are Schools 1–8 graduates, 4,767 individuals (76.2%) are Imperial University graduates, and 982 individuals (15.7%) are top-ranking officials. The share of Imperial University graduates is greater than that of Schools 1–8 graduates because the number of higher schools increased from 8 to 25 in 1919–1930. Among 982 top-ranking officials, 701 officials (71.4%) are Schools 1–8 graduates, and 891 officials (90.7%) are Imperial University graduates. The HCSE consisted of preliminary and main exams, and because Imperial University law graduates were the only group exempted from the administrative preliminary exams from 1894 to 1922, they had substantial advantages in passing the HCSE (Hata, 1981, pp.663-666; Spaulding, 1967, Chapter 12). This exemption, however, does not explain why Imperial University graduates had such a high share in the top-ranking officials.
@@ A.4.3 Variation in Years Taken to be Promoted to Top-ranking Positions
One potential threat to our identification in Table 4 Panel (b) is a possibility that the number of available top-ranking positions happened to have increased during the periods of centralization. However, we argue that even if this was the case, it is not likely to affect our results, since our cohort is defined by the year of entering a higher school or its equivalent, not by the year of becoming top-ranking officials. Namely, as long as individuals in a given cohort were not promoted to a top-ranking position in the same year, a potential correlation between the number of top-ranking positions and the lagged periods of centralized admissions does not bias our results.
To examine this possibility, we first randomly selected two cohorts exposed to centralization (1903 and 1916) and two cohorts exposed to decentralization (1913 and 1922). For all top-ranking officials who graduated from Schools 1–8 in these cohorts, we searched the years in which these officials were appointed to their first top-ranking positions. Using online searches to find biographical information for each official, we obtained necessary information for 82% of these officials. As shown in Appendix Figure A.9, within each cohort, the number of years taken from entering a higher school to the appointment for the first top-ranking position varied widely from 20 to 30 years. This within-cohort variation is a sum of two variations, the first variation in the number of years taken from entering a higher school to passing the exams and the second variation in the number of years taken from passing the exams to becoming a top-ranking official. Therefore, it is unlikely that the result in Table 4 Panel (b) is driven by a greater number of available top-ranking positions that coincided with the periods of centralization.
@@ A.4.4 Connections and Signaling Hypotheses
We investigate whether connections or signaling was an important mechanism through which meritocratic centralization increased the total number of top-ranking officials. First, to explore the connection channel, we reexamine the result of Table 4 Panel (b) Column (3) by controlling for the number of exam passers from the same schools (Schools 2–8) in the same cohort, instead of the total number of exam passers. This variable takes into account the size of same-cohort colleagues from the same schools who are likely to have network values. The coefficient of centralization remains sizable (2.72) and statistically significant at the 1% level, implying a 24% increase in the number of top-ranking officials from these schools relative to the mean during decentralization (11.22). Second, to investigate the signaling channel, we further restrict our sample to top-ranking officials who graduated from both Schools 2–8 and Tokyo Imperial University, the most prestigious institution with the strongest signaling value. We regress the number of this subgroup of top-ranking officials on centralization, controlling for the number of exam passers from the same schools in the same cohort. If signaling is a major channel, we expect the effect of centralization among this subgroup to be larger. The coefficient of centralization decreases to 1.76, implying a 18% increase in the number of top-ranking officials from these schools relative to the mean during decentralization (10.00). This effect size is smaller than the effect size of 24% above.
@@
Notes: This figure summarizes each country and territory’s college admission system today. Dark red color (e.g. Norway): Regionally- or nationally-centralized college admissions where a single-application, single-offer assignment algorithm (well-defined rule) is used to make admissions to both public and private universities. Medium orange color (e.g. Brazil): Semi-centralized college admissions defined as either (1) there is a centralized system, but not all universities (such as private universities) are included in the single-application, single-offer system or (2) students submit a single application and receive multiple offers. Light orange color (e.g. U.S.): Decentralized college admissions where each college defines its own admission standards and rules. Yellow with diagonal lines (e.g. Chad): Not enough information available or if the country or territory does not have tertiary institutions. We summarize the information sources at https://www.scribd.com/document/437545135/Online-Appendix191018. See Section 2 for discussions about this figure.
@@
Notes: This figure shows the locations of Schools 1–8 and the location of the Tokyo area (in dark gray color) defined as a set of 7 prefectures that are within 100 km from Tokyo (i.e., Tokyo, Chiba, Kanagawa, Saitama, Ibaraki, Tochigi, and Gunma prefectures). See Sections 2 and 3.5 for discussions about this figure.
@@
Notes: This table shows changes in the admission system of National Higher Schools (including Schools 1–8) from 1900 to 1930. In 1901, all schools held their exams on the same date, with the exception of newly established School 7. In 1908, all schools held their exams on the same date, with the exception of Schools 7 and 8. See Moriguchi (2021) for historical details and the sources of information. See Section 2.1 and Appendix Section A.1 for discussions about this table.
@@
Notes: This figure is a reprint of the assignment algorithm of the centralized admission system stated in the Ordinance of the Ministry of Education No.4 published in Government Gazette No.1419, pp.580-581, on April 27, 1917. See Sections 2 and 3.4 for an English translation and discussions.
@@
No. of middle-school graduates 533.35 617.82 359 1457
@@
No. of corporate executives with a positive amount of tax payment 8.44 9.57 6 705
@@ B. Application Distance
Notes: This table uses the prefecture-year level data in 1900–1930. The dependent variable is the number of students from birth prefecture $p$ who entered one of Schools 1–8 in year $t$. “Population in prefecture” is population in prefecture $p$ in year $t$. “GDP per capita in prefecture” is real gross value-added per capita in prefecture $p$ in year $t$. “Middle-school graduates in prefecture” is the number of students who graduated from middle schools in prefecture $p$ in year $t$. “Middle-school graduates in nearby prefectures” is the number of students who graduated from middle schools in the prefectures within 100 km from prefecture $p$ (excluding prefecture $p$) in year $t$. “Share of applicants to School 1 (under decentralization)” is the share of applicants to School 1 among all applicants to Schools 1–8 in prefecture $p$ under the decentralized system in 1916 (the only year for which the data is available). All variables interacted with “Centralized” are standardized to be mean 0 and standard deviation 1. We control for year fixed effects, prefecture fixed effects, and the number of higher schools other than Schools 1–8 in prefecture $p$ in year $t$. We also control for “Born in school’s prefecture,” “Born near school’s prefecture (1–100 km),” and “Born near school’s prefecture (100-300 km)” as in Table 1. Parentheses contain p-values based on standard errors clustered at the prefecture level. Square brackets contain p-values based on standard errors clustered at the year level. ***, **, and * mean significance at the 1%, 5%, and 10% levels, respectively. See Section 3.5 for discussions about this table.
@@
No. of middle-school graduates 0.00493 -0.000942 -0.150 0.00134 0.0656 -0.429 0.00194 1.448 0.00150
@@
Mean dep var 23.41 2.007 0.180 10.78 0.217 19.03 13.56 0.291 228.4 0.174
Notes: Columns 1–7 test if important institutional variables are correlated with the timing of centralization using year-level data. Columns 8–10 do the same our main short-run outcomes using year-level data. All numbers are at the national-level from 1900 to 1930. The numbers of middle school graduates, entrants, and applicants are denominated by 1,000. In all regressions, quadratic time trends (i.e. trend and trend squared, where the trend is defined by “year - 1899”) are controlled. Parentheses contain p-values based on Newey-West standard errors with the maximum lag order of 3. See Section 3.6 for discussions about this table.
Table A.8: Occupations of Individuals Listed in the PIR (1939)
(1) All individuals listed (2) Sampled cohorts (3) Top 0.05% income earners (4) Medal recipients
Corporate executives and managers 0.660 0.672 0.803 0.320
Politicians and bureaucrats 0.075 0.070 0.030 0.213
Scholars, lawyers, and artists 0.119 0.142 0.048 0.473
Engineers and physicians 0.090 0.110 0.048 0.259
Military personnels 0.045 0.047 0.019 0.000
Landlords 0.048 0.033 0.060 0.001
Imperial and peerage family members 0.006 0.004 0.004 0.016
None of the above 0.130 0.109 0.098 0.087
No. observations 55,742 28,423 3,864 4,739
Notes: This table shows occupational distributions of individuals listed in the PIR (1939). Column (1) shows the share of each specified occupational category among all individuals listed in the PIR (1939). Column (2) restricts the sample to cohorts who were born in 1880–1894. Columns (3) and (4) further restrict the sample to the top 0.05% income earners and civilian medal recipients (as defined in Section 4.1), respectively. Occupational categories are not mutually exclusive, except for “None of the above” which is defined as individuals not included in the specified categories. See Section 4.1 for discussions about this table.
Table A.9: Definitions of Occupational Categories of Elites
Notes: This table provides the precise definitions of occupational categories of elites. Individuals in each category are defined as a subset of individuals listed in the PIR. See Section 4.1 for the analysis using these categories.
@@
Notes: This figure plots the sampling rate of the high income earners listed in PIR (1939) by the income level expressed as a top percentile of the national income distribution. The sampling rates and the top income percentiles are computed from the complete count data in the National Tax Bureau Yearbook. The vertical lines indicate the top 0.05% and 0.01% thresholds used in our analysis. See Section 4.1 for discussions about this figure.
@@
Notes: These figures compare the number of high income earners in each prefecture listed in the PIR (1939) and the complete count of high income earners in each prefecture reported in the National Tax Bureau Yearbook (1936). The vertical axis is the log of the number of individuals in the PIR who earned more than 50,000 JPY taxable income (corresponding to the top 0.01% income group) or 16,950 JPY taxable income (corresponding to the top 0.05% income group) in 1938. The horizontal axis is the log of the number of individuals in tax statistics who earned more than 30,000 JPY taxable income (corresponding to the top 0.013% income group) or 10,000 JPY taxable income (corresponding to the top 0.08% income group) in 1936 (the closest year to 1938 for which prefecture-level tax statistics are available). See Section 4.1 for discussions about these figures.
@@
Figure A.7: Long-run Impacts of Centralization: Geographical Origins of Other Elites
Notes: This figure shows additional difference-in-differences plots that compare the average number of elites born in prefectures inside and outside the Tokyo area by cohort (see Figure 3). The plots are based on the data from PIR (1939), which covers cohorts who were born in 1879–1894 and turned age 17 (main application age) in 1896–1911. The vertical axis shows the average number of indicated elites (per prefecture) born in the indicated area in the indicated cohort. The cohorts are colored according to their intensity of exposure to the centralized admissions in 1902–07, where the darker color indicates the higher intensity. See Section 4.1 for discussions about this figure.
Table A.11: Long-run Impacts: Other Elites
@@
Age 17 under centralization 1.36 0.66 0.46 0.18
$\times$ Tokyo area ($<$100 km) (0.004)*** (0.048)** (0.038)** (0.420)
[0.003]*** [0.125] [0.091]* [0.544]
@@
Mean dep var 4.29 2.46 1.92 1.11
Mean dep var (Tokyo area 5.41 3.13 2.52 1.16
@@
Imperial Top 0.01% Top 0.05% Medal Corporate Top Imperial All
Univ. income income recipients executives politicians & Univ. occupational
grads earners earners bureaucrats professors elites
Panel A: Baseline Specification
Age 17 under centralization 3.05 0.68 1.65 2.61 1.94 0.68 0.45 9.27
    $\times$ Tokyo area ($<$100 km) (0.003)*** (0.007)*** (0.048)** (0.008)*** (0.020)** (0.019)** (0.067)* (0.068)*
[0.000]*** [0.038]** [0.019]** [0.000]*** [0.081]* [0.022]** [0.193] [0.001]***
Panel B: Adding Control Variables
Age 17 under centralization 1.70 0.58 1.53 2.18 1.61 0.45 0.41 6.48
    $\times$ Tokyo area ($<$100 km) (0.021)** (0.013)** (0.007)*** (0.001)*** (0.007)*** (0.028)** (0.095)* (0.007)***
[0.019]** [0.010]** [0.013]** [0.001]*** [0.081]* [0.095]* [0.207] [0.009]***
Observations 611 611 611 611 611 611 611 611
Cohort FE, Birth pref. FE Yes Yes Yes Yes Yes Yes Yes Yes
Mean dep var 8.68 1.23 5.16 6.08 8.41 1.32 0.83 37.72
Mean dep var (Tokyo area 10.38 1.41 6.57 6.61 11.39 1.18 1.02 49.43
@@
Imperial Top 0.01% Top 0.05% Medal Corporate Top Imperial All
Univ. income income recipients executives politicians & Univ. occupational
grads earners earners bureaucrats professors elites